(x - 2)(x - 2i)(x + 2)(x - 2i) is equal to
1
Expert's answer
2021-03-31T13:46:53-0400
"(x - 2)(x - 2i)(x + 2)(x - 2i)"

"=(x^2-4)(x^2-4i-4)"

"=x^4-4x^2i-4x^2-4x^2+16i+16"


"=x^4-8x^2 +16-(4x^2-16)i"